• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 08:22:21 on November26,2008

    Entry number 250868

    keyword=Owl shift summary

    SL : A. Camsonne
    TO : Y. Zhang
    3rd : M. Iodice


    0:00 taking over from swing shift

    ~0:15 Problem with EPICS solved by asking to reboot the Hall A gateway

    00:30 moved to BeO target

    Beam seems still a bit offset.
    Simon request a shift by -1 mm
    currennt position is
    BPM A ( 0.5,1.7)
    BPM B ( -1,1.0 )

    asked for
    BPM A ( 0.,1.5)
    BPM B ( -1.5,1.1 )

    00:41 looks well centered on BeO

    00:48 Target moved to reference cell asked for 2uA rastered beam

    1:12 Rates do not change so we will stay at this postion

    BPM A ( 0.,1.5)
    BPM B ( -1.5,1.1 )


    1:14 Run empty reference cell 5 uA

    1:38 Filled reference cell with N2

    2:31 ROC9 rebooted

    2:58 ROC rebooted by itself 3 times... use remote reset and was fine after

    3:39 Switched to H2 110 psig in reference cell

    6:34 Switched to He3 reference cell 110 psig

    6:50 End of run failed and Happex DAQ crashed

    7:35 HV alarm, Beamline and HRS red and display not updating but HV looks fine on portserver

    8:04 HV beamline tripped.. turned it off and on remotely
    handing data taking to day shift

    ---------------------------
    Took few compton runs with low electron rate in garage position

    Did RICH Track and Hold scan it seems we can gain about 10% of signal by reducing the delay on the T/H. Will do at the next access. We will run



    --------------------------
    BeO run :

    run 3924 No raster

    Reference cell empty runs for alignement checks :
    Run 3925 3x4 mm MCC units
    Run 3926 4x5.5 mm MCC units
    Run 3927 5x6.5 mm MCC units


    Reference cell empty run :
    Run 3928 raster 3x4 mm MCC gives 3x3 mm at target , short run while setting prescales
    Run 3929 run empty reference cell 5uA

    Reference cell run N2 26 psig
    Run 3630 running while filling the cell
    Run 3631 12 uA BigBite HV being changed
    Run 3632 12 uA 400 Kevents
    Run 3633 junk
    Run 3634 12 uA Rich T/H -120 ns
    Run 3635-3938 Junk run ROC9 problem
    Run 3939 12 uA RICH T/H +120 ns
    Reference cell H2 110 psig
    Run 3940 12uA RICH +64ns 1 Mevents
    Run 3941 12uA RICH -64ns 1 Mevents
    Run 3942 12uA RICH -64ns 0.65 Mevents RICH at 0 ns

    Reference cell He3 110 psig
    Run 3943 12 uA short run while changing prescales
    Run 3945 12 uA 0.8 M
    Run 3946 12 uA running